# The Merging Galaxy Cluster Environment Affects the Morphology of Radio-AGN

### v1

@@ 3.6 Statistical Methods
We want to compare the proportion of observed radio morphology classes in mergers and non-mergers within the uncertainties. To do this, we express our results as proportions (), where is the number of radio-AGN of a particular radio morphology in mergers or non-mergers and is the total number of radio sources for the respective dynamical state.
Given the small number statistics and asymmetric errors on the proportions measured, care is necessary in quoting and defining uncertainties. We take a Bayesian approach and give credible intervals on the true proportion given the observed and . For a Jeffreys prior, the posterior probability distribution of is a Beta distribution:
and hence the credible interval can be taken to be the relevant percentile of the Beta distribution (e.g. 16 and 84% for a range as used here). It is conventional to set the lower range to 0 if and the upper range to 1 if (the Jeffreys interval, which is indicated as the lower and upper uncertainties on Figure 3), as implemented in the astropy.stats function binom_conf_interval.
It is important to note that, given the asymmetric nature of these errors and the fact that the posterior distribution about is far from being a Gaussian, it is not possible simply to combine the quoted ‘errors’ (credible interval) in order to compare two different proportions with one another, or to multiply by a constant to get a confidence range. If we observe two proportions and where , then the Bayesian way of asking the question ‘is significantly greater than ’ is ‘is there a high posterior probability that ?’ This could be found exactly by convolving the two posterior distribution functions, but for simplicity we simply subtract samples drawn from and , which allows us to estimate this posterior probability to adequate precision (listed as Pr(MNM) in Table 2). We map the Pr(MNM) to a statistical significance level using a one-tailed test by inverting the cumulative distribution function of a Gaussian. Although an approach framed as ‘is there a difference?’ or in other words testing if Pr(MNM) or Pr(MNM), may seem less biased, there are cases where the proportion is 0, making it unreasonable to test for any direction other than greater. For consistency, we maintain a one-tailed approach throughout the analysis. These statistical significances are quoted and interpreted in Section 4.1, taking the standard approach that 68%, 95%, and 99.7% correspond to 1, 2, and 3, respectively.
@@ 4 Results
After determining consensus and radio-AGN delineation, we analyze the trends in our data based on morphology (see §4.1) and AGN classification (see §4.2).
@@ 4.1 Combined Classifications and Results
Compact Extended: This group of classifications includes only compact extended (CE) sources. We keep CE separate from the compact objects as they exhibit a single extension, but differ from clearly extended, as they have a clear compact core.
@@
Jetted: This group combines sources that have FR I, FR II, or FRM classifications. First, we combine these classes in order to increase the number of sources in this class and second, because they both display unbent jets that have clear two-sided jet morphology. It it not important to distinguish between FRI and FRII for this work because they both have unbent, two-sided jets which are similar morphologically compared to the other types of morphologies considered in this work (e.g., compact, bent, etc.)
@@
Disturbed: This group of classifications combines Bent Tail (BT) and Head Tail (HT) sources. Tailed sources are indicative of ram pressure stripping as a galaxy navigates the dense ICM. The degree to which a source is bent is beyond the scope of this work, thus we simply combine all sources that have jets that are bent to some degree to differentiate the bent from unbent radio morphologies.
@@
Enumerated in Table 2 and visually in Figure 3, we report the proportion and uncertainties (described in §3.6) of the combined morphology classifications. In non-mergers, we observe a population proportions of compared to in mergers. Similarly in the CE category, in non-mergers, the value is compared to in mergers. Using the posterior probabilities of the C and CE across mergers and non-mergers, the merging population does not exhibit statistically significantly greater proportions of either C and CE sources compared to non-mergers (in other words 1). The calculation follows from mapping Pr(MNM), where Pr(MNM) is equivalent to (M NM 0) divided by N samples, to the corresponding values discussed in Section 3.6.
@@
Interestingly, we find extended (E) sources only in non-mergers. In non-mergers, there is a population proportion of compared to in mergers. Mergers do not display a significantly greater () population proportion of E sources as compared to the non-mergers. Sources classified as E range greatly in morphology. This classification aimed to describe all extended, amorphous emissions that could not be unambiguously classified as compact or jetted. Based on a preliminary, quick optical counterpart examinations by eye, E sources may arise from a variety of physical mechanisms including but not limited to starburst face-on spirals, jellyfish galaxies, and cluster member mergers. The physical reasoning for heightened E sources in non-mergers is unclear, and a more detailed follow-up with optical counterparts would pose an interesting future direction. The origin of the emission mechanism for E sources is beyond the scope of this paper and requires a multi-wavelength approach for robust analysis.
@@
We find that the proportion of jetted sources in mergers is not significantly (1) greater than the proportion of jetted sources in non-mergers. We find the observed population proportion to be for mergers and non-mergers to be . The jetted sources encompass all clear FR-type morphologies.
@@
Lastly, we find that the merging systems have a greater proportion of disturbed sources than non-mergers, within the error. We find that mergers and non-mergers have a population proportion of and , respectively. Mergers observe a greater proportion of disturbed sources as compared to non-mergers, at the 3 level. In other words, we find the posterior probability that Pr(MNM) is greater than the 3 equivalent probability of 0.997. After visually re-examining the images of the disturbed sources, we find that they appear to be both bent and head tails which are jetted sources that have been affected by ram pressure () causing the jets to become bent to some degree (Miley et al., 1972; Owen & Rudnick, 1976; Begelman et al., 1979; O’Dea & Owen, 1985).
@@
In summary, we find mergers do not exhibit greater proportions of C sources, CE sources, and Jetted sources while they contain no E sources and significantly more disturbed sources when compared to the non-mergers. Non-mergers have comparable proportions of C sources, CE sources, Jetted sources, and contain all the E sources and significantly fewer disturbed sources as compared to the mergers. We believe that this is a result of the large-scale bulk motions of the ICM that are present in the merging environment (see discussion in §5).
@@ 4.2 Comparing Radio and Optical AGN
In Table 3, we report the statistics of radio-AGNs that are (a) also identified as an optical AGN (AGNRad-Opt) and (b) only identified in the radio (AGNRad) across both dynamical states. For the AGNRad-Opt, we cross-reference our radio-AGN sources with the Bilton et al. (2020) optical AGN catalog and the sources that are in both samples are counted as AGNRad-Opt. Conversely, the radio-AGN in our sample that are not identified by Bilton et al. (2020) as an optical AGN and are classified as radio-AGN only and are thus counted as AGNRad.
In comparing the proportions of AGNRad-Opt and AGNRad in mergers and non-mergers, we only find one marginally significant result: there are more AGNRad than AGNRad-Opt in merging systems taking into account the errors at a 2 level. In other words, the posterior probability that Pr(AGNRad AGNRad-Opt) within mergers is 0.96. Less significant but noteworthy, we find mergers exhibit a marginally greater proportion of AGNRad at the 1.5 level as compared to non-mergers, or the posterior probability that Pr(MNM) is 0.86. Next, comparison of AGNRad and AGNRad-Opt proportions within non-mergers yield a Pr(AGNRadAGNRad-Opt) equivalent statistical significance less than one, where the Pr(AGNRad AGNRad-Opt) is 0.59. Lastly, the comparison of AGNRad and AGNRad-Opt across mergers and non-mergers, yield a Pr(MNM) equivalent statistical significance of less than one, where Pr(MNM) is 0.14.
@@ 5 Discussion
This paper examines the population proportion of radio-AGN morphologies and radio-AGN and optical AGN optical across merging and non-merging clusters. Our findings suggest that mergers have a statistically significant, at a 3 level, increased prevalence of disturbed radio sources and that mergers and non-mergers have a relatively equal prevalence of compact and jetted sources. Notably, a merging cluster environment enables large scale bulk motion in the ICM, increases the potential for more galaxy-galaxy interactions, and heightens the overall disorder in the system. A dynamic environment such as that found in merging systems creates more opportunities for radio jets to interact with their surrounding and become bent (Morris et al., 2022; O’Dea & Baum, 2023) or “disturbed”.
@@
It has already been suggested in the literature that merging cluster systems can affect radio-AGN morphology. There are historically two types of bent-tail radio-AGN: WATs (larger opening angle between the two radio jets) and Narrow Angle Tails (NATs, which have a smaller opening angle between the two radio jets). WATs seem to be fundamentally different from Narrow Angle Tails (NATs) in a variety of ways. The main difference is that WATs are cluster-center objects near the bottom of the cluster potential, and thus WAT host galaxies must be nearly at rest (Owen & Rudnick, 1976; Quintana & Lawrie, 1982; Eilek et al., 1984). Therefore, bent WAT morphology cannot be explained by ram pressure due to the motion of the host galaxy. Instead, the WAT bent morphology has been posited to be a result of merging cluster-cluster systems (Roettiger et al., 1993; Pinkney et al., 1994; Burns, 1998; Sakelliou & Merrifield, 2000; Burns et al., 2002) where the bulk scale motions of the ICM are responsible for the ram pressure that bends the tails. And even though there is still debate of the origin of Hybrid Morphology Radio Sources (intrinsic or environmental), an asymmetric environment has been posited as one origin (Gopal-Krishna & Wiita, 2000; Kapińska et al., 2017) which could be created in a cluster-cluster merger (Gawroński et al., 2006). Therefore, it is not unexpected that a merging system could cause a change in morphology of jetted systems as we see in this work.
@@
We note that we only see head tails in merging systems (see Table 6). Head tails are extreme versions of bent tail systems and are a product of stronger ram pressure either from a denser medium or a higher velocity of the source (; Miley et al. 1972; Terni de Gregory et al. 2017). Because only the most disturbed systems (HTs) are found in merging systems, this may indicate that merging systems allow for more opportunities for extreme ram pressure. An interesting follow-up would be to investigate the placement of these head tail systems with respect to the cluster center.
@@
Our findings suggest that mergers have a maginally significant increased prevalence of radio-AGN that are also identified as an optical AGN. If we assume that AGNRad can be roughly equated to an inefficiently accreting engine or low-excitation radio galaxies (LERGs) and the AGNRad-Opt can be roughly equated to an efficiently accreting engine or high-excitation radio galaxies (HERGs) (Hardcastle et al., 2009; Buttiglione et al., 2010; Best & Heckman, 2012) then we see that there are more LERGs than HERGs in merging clusters. Though we do not do a full HERG/LERG classification in this analysis (this was not the goal of this work so potential biases exist such as the fact that this sample was not created to have a S/N target required for detecting the spectral lines needed to classify HERG/LERGs are present), we note that it would be interesting future work. In fact, the environment affecting the incidence of LERGs versus HERGs is not unprecedented. A few early works find that in the local universe LERGs occupy more dense environments than HERGs (Prestage & Peacock, 1988; Allington-Smith et al., 1993; Zirbel, 1997; Best, 2004), which was confirmed by Miraghaei & Best (2017). On the other hand, Massaro et al. (2019, 2020) do not find a difference in HERG/LERG environments in the local universe and note that the difference in environments of HERGs and LERGs seen in other works could be due to selection bias. However, when Ineson et al. (2015) compare the large-scale environments of HERG and LERG in the local universe () to that of those at , they do find a difference and note that there is strong evidence of a difference between the large-scale environments of HERGs and LERGs.
@@ 5.1 Discussion of Potential Sources of Uncertainty
There are several sources of error which we discuss here. An inherent source of error exists in the limitations of our classifications. Each source presents a unique radio emission profile, and while we aimed to encompass the bulk trends with the classification schemes, some intricacies may have eluded our classification process. Our morphological classification system does not encompass every type of radio galaxy morphology and was not designed to. We acknowledge the presence of jellyfish galaxies in our sample. Notably, we find examples of WC sources exhibiting faint extended emission that do not align exactly with the CE as defined here in this work, but are noteworthy. In fact, two WC sources in Abell 2255 are found in the existing literature to be ram-pressure stripped galaxies (Ignesti et al., 2023). Further we cross-match our sample with the LoTSS jellyfish catalog from Roberts et al. (2021) and we find that four of our radio AGN are identified as jellyfish galaxies. We find that the jellyfish galaxies are equally spread between the WC, SC, CE, and E categories. However we note that the LoTSS jellyfish catalog Roberts et al. (2021) only includes data from LoTSS DR2 data and up to half clusters in this work have LoTSS data that have been taken since DR2. Similarly, we identified sources displaying FR-type morphologies that do not precisely fit FRM or FRI/II categories, potentially due to image artifacts, yet they exhibit distinctive and intriguing morphologies. To maintain the scope of this project, we limit our results to only the original Zooniverse classification results.
@@
An additional source of error that arises inherently in our classification-based project is disagreement. First there are only 71 sources with complete agreement from all classifiers (consensus level 5) and the remaining 90 sources do not have complete agreement. On the other end of the spectrum within our data set, there were 26 sources that did not have a majority agreement (consensus level 2). Involving multiple classifiers was a necessary step to obtain robust classifications, even though it introduces some margin for error. There is a growing movement to move towards having multiple classifications or “tags” per source in the radio classification community in order to better describe the complex nature of the morphology of radio galaxies (Rudnick, 2021). For this exploratory work, we chose to have only one classification per object but that leaves room for “error” in that we may not capture the full, complex morphology of the source.
@@
Another source of error is that of the dynamical state classification of the host cluster. It is difficult to obtain a binary classification of the dynamical state of clusters into “merging” and “relaxed.” Galaxy clusters are a direct result of hierarchical merging systems evolving over cosmic time and the clusters we examine are snapshots in time, each characterized by varying degrees of merging activity. There is an ongoing debate surrounding the optimal approach for determining the dynamical state of a galaxy cluster. Our study is based on the Bilton et al. (2020) data set which classifies clusters as merging or non-merging based on a kinematic analysis. However existing literature suggests certain clusters that were classified as a ‘non-merging’ cluster by Bilton et al. (2020) may in fact be undergoing merging processes based on X-ray observations which trace ICM turbulence (Nulsen et al., 2013; Blanton et al., 2011; Bailey et al., 2016; Drabent et al., 2015; Yu et al., 2016; Wen & Han, 2013). Another way to classify the dynamical state of a cluster is based on the presence of central radio halos and radio relics which can serve as an effective indicator of ongoing merger events. Notably, at least two clusters (Abell 1367 and Abell 1656) in this sample, classified as non-merging using the Bilton et al. (2020) method, are identified as mergers based on radio halo observations (Kim et al., 1990; Gavazzi, 1978; Bonafede et al., 2022; Ge et al., 2019; Zhang et al., 2023).
@@
With these facts in mind, the clusters that Bilton et al. (2020) identified as merging and non-merging via kinematics might not have that classification if one were to use a different type of analysis. Though it is difficult to get uniform, multi-wavelength data for a large sample of clusters, and depending on the science goals, a multi-wavelength approach would be the most robust best way to determine the dynamical state of a cluster.
@@
A last source of error is the choice to use a global RMS for the cutout contours. The choice of a global RMS per cluster, inherently, may lead to the loss of smaller-scale emission for individual sources where the RMS may be different from the global RMS. Given our original dataset from the Bilton et al. (2020) catalog, which comprises over 1000 sources, we were confronted with a trade-off between efficiency and sensitivity. Consequently, emission of lower surface brightness, such as that associated with ram pressure stripping, may not be fully captured. Existing work suggests at least two WC sources from Abell 2255 are ram-pressure stripped galaxies (Ignesti et al., 2023). However, in §3.1 we investigate the effect of choosing global versus a local RMS and find that our choice of a global, cluster RMS does not affect our results.

### v3

@@ 3.6 Statistical Methods
We want to compare the proportion of observed radio morphology classes in mergers and nonmergers within the uncertainties. To do this, we express our results as proportions (/), where is the number of radio AGN of a particular radio morphology in mergers or nonmergers and is the total number of radio sources for the respective dynamical state.
Given the small number statistics and asymmetric errors on the proportions measured, care is necessary in quoting and defining uncertainties. We take a Bayesian approach and give credible intervals on the true proportion given the observed and . For a Jeffreys prior, the posterior probability distribution of is a Beta distribution,
and hence the credible interval can be taken to be the relevant percentile of the Beta distribution (e.g. 16% and 84% for a range as used here). It is conventional to set the lower range to 0 if and the upper range to 1 if (the Jeffreys interval, which is indicated as the lower and upper uncertainties in Figure 3), as implemented in the astropy.stats function binom_conf_interval.
It is important to note that, given the asymmetric nature of these errors and the fact that the posterior distribution about / is far from being a Gaussian, it is not possible simply to combine the quoted ‘errors’ (credible interval) in order to compare two different proportions with one another, or to multiply by a constant to get a confidence range. If we observe two proportions and where //, then the Bayesian way of asking the question ‘Is significantly greater than ’ is, ‘Is there a high posterior probability that ?’ This could be found exactly by convolving the two posterior distribution functions, but for simplicity we simply subtract samples drawn from and , which allows us to estimate this posterior probability to adequate precision (listed as Pr(MNM) in Table 2). We map the Pr(MNM) to a statistical significance level using a one-tailed test by inverting the cumulative distribution function of a Gaussian. Although an approach framed as ‘Is there a difference?’ or in other words testing whether Pr(MNM) or Pr(MNM), may seem less biased, there are cases where the proportion is 0, making it unreasonable to test for any direction other than greater. For consistency, we maintain a one-tailed approach throughout the analysis. These statistical significances are quoted and interpreted in Section 4.1, taking the standard approach that 68%, 95%, and 99.7% correspond to 1, 2, and 3, respectively.
@@ 4 Results
After determining consensus and radio AGN delineation, we analyze the trends in our data based on morphology (see §4.1) and AGN classification (see §4.2).
@@ 4.1 Combined Classifications and Results
Compact extended: This group of classifications includes only compact extended (CE) sources. We keep CE separate from the compact objects, as they exhibit a single extension but differ from clearly extended, as they have a clear compact core.
@@
Jetted: This group combines sources that have FR I, FR II, or FRM classifications. We combine these classes, first in order to increase the number of sources in this class, and second because they both display unbent jets that have clear two-sided jet morphology. It it not important to distinguish between FR I and FR II for this work because they both have unbent, two-sided jets that are similar morphologically compared to the other types of morphologies considered in this work (e.g., compact, bent, etc.)
@@
Disturbed: This group of classifications combines bent-tail (BT) and head-tail (HT) sources. Tailed sources are indicative of ram pressure stripping as a galaxy navigates the dense ICM. The degree to which a source is bent is beyond the scope of this work; thus, we simply combine all sources that have jets that are bent to some degree to differentiate the bent from unbent radio morphologies.
@@
Enumerated in Table 2 and visually in Figure 3, we report the proportion and uncertainties (described in §3.6) of the combined morphology classifications. For the compact category of sources, in nonmergers, we observe population proportions of compared to in mergers. Similarly in the CE category, in nonmergers, the value is compared to in mergers. Using the posterior probabilities of the C and CE across mergers and nonmergers, the merging population does not exhibit statistically significantly greater proportions of either C or CE sources compared to nonmergers. In other words, for sources in the C and CE categories their corresponding Pr(MNM) are 0.325 and 0.146, respectively, which both map to the equivalent of . The calculation follows from mapping Pr(MNM), where Pr(MNM) is equivalent to (M NM 0) divided by N samples, to the corresponding values discussed in Section 3.6.
@@
Interestingly, we find E sources only in nonmergers. In nonmergers, there is a population proportion of compared to in mergers. Mergers do not display a significantly greater () population proportion of E sources as compared to the nonmergers. Sources classified as E range greatly in morphology. This classification aimed to describe all extended, amorphous emissions that could not be unambiguously classified as compact or jetted. Based on a preliminary, quick optical counterpart examination by eye, E sources may arise from a variety of physical mechanisms, including but not limited to starburst face-on spirals, jellyfish galaxies, and cluster member mergers. The physical reasoning for heightened E sources in nonmergers is unclear, and a more detailed follow-up with optical counterparts would pose an interesting future direction. The origin of the emission mechanism for E sources is beyond the scope of this paper and requires a multiwavelength approach for robust analysis.
@@
For the jetted category of sources (the jetted sources encompass all clear FR-type morphologies), we find the observed population proportion for mergers is and that for nonmergers to be . And using the posterior probabilities of the jetted sources across mergers and nonmergers, we find that the merging population does not exhibit a statistically significant greater proportion of jetted sources (Pr(MNM) of 0.68 which is the equivalent of ).
@@
Lastly, we find that the merging systems have a greater proportion of disturbed sources than nonmergers, within the error. We find that mergers and nonmergers have a population proportion of and , respectively. And using the posterior probabilities of the disturbed sources across mergers and nonmergers, we find that the merging population does exhibit a statistically significant greater proportion of disturbed sources as compared to nonmergers (Pr(MNM) of 0.997 which is equivalent to 3). After visually re-examining the images of the disturbed sources, we find that they appear to be both bent and head-tails, which are jetted sources that have been affected by ram pressure (), causing the jets to become bent to some degree (Miley et al., 1972; Owen & Rudnick, 1976; Begelman et al., 1979; O’Dea & Owen, 1985).
@@
In summary, we find that mergers do not exhibit greater proportions of C sources, CE sources, and jetted sources while they contain no E sources and significantly more disturbed sources when compared to the nonmergers. Nonmergers have comparable proportions of C sources, CE sources, and jetted sources and contain all the E sources and significantly fewer disturbed sources as compared to the mergers. We believe that this is a result of the large-scale bulk motions of the ICM that are present in the merging environment (see discussion in §5).
@@ 4.2 Comparing Radio and Optical AGN
In Table 3, we report the statistics of radio AGNs that are (a) also identified as an optical AGN (AGNRad-Opt) and (b) only identified in the radio (AGNRad) across both dynamical states. For the AGNRad-Opt, we cross-reference our radio AGN sources with the Bilton et al. (2020) optical AGN catalog, and the sources that are in both samples are counted as AGNRad-Opt. Conversely, the radio AGN in our sample that are not identified by Bilton et al. (2020) as an optical AGN and are classified as radio AGN only and are thus counted as AGNRad.
In comparing the proportions of AGNRad-Opt and AGNRad in mergers and nonmergers, we only find one marginally significant result: there are more AGNRad than AGNRad-Opt in merging systems taking into account the errors at a 2 level. In other words, the posterior probability that Pr(AGNRad AGNRad-Opt) within mergers is 0.96. Less significant but noteworthy, we find that mergers exhibit a marginally greater proportion of AGNRad at the 1.5 level as compared to nonmergers, or the posterior probability that Pr(MNM) is 0.86. Next, comparison of AGNRad and AGNRad-Opt proportions within nonmergers yields a Pr(AGNRadAGNRad-Opt) equivalent statistical significance less than one, where the Pr(AGNRad AGNRad-Opt) is 0.59. Lastly, the comparison of AGNRad and AGNRad-Opt across mergers and nonmergers, yields a Pr(MNM) equivalent statistical significance of less than 1, where Pr(MNM) is 0.14.
@@ 5 Discussion
This paper examines the population proportion of radio AGN morphologies and radio AGN and optical AGN across merging and nonmerging clusters. Our findings suggest that mergers have a statistically significant (at a 3 level) increased prevalence of disturbed radio sources and that mergers and nonmergers have a relatively equal prevalence of compact and jetted sources. Notably, a merging cluster environment enables large-scale bulk motion in the ICM, increases the potential for more galaxygalaxy interactions, and heightens the overall disorder in the system. A dynamic environment such as that found in merging systems creates more opportunities for radio jets to interact with their surroundings and become bent (Morris et al., 2022; O’Dea & Baum, 2023) or “disturbed.”
@@
It has already been suggested in the literature that merging cluster systems can affect radio AGN morphology. There are historically two types of bent-tail radio AGN: wide-angle tails (WATs; which have a larger opening angle between the two radio jets) and narrow-angle tails (NATs; which have a smaller opening angle between the two radio jets). WATs seem to be fundamentally different from NATS in a variety of ways. The main difference is that WATs are cluster-center objects near the bottom of the cluster potential, and thus WAT host galaxies must be nearly at rest (Owen & Rudnick, 1976; Quintana & Lawrie, 1982; Eilek et al., 1984). Therefore, bent WAT morphology cannot be explained by ram pressure owing to the motion of the host galaxy. Instead, the WAT bent morphology has been posited to be a result of merging clustercluster systems (Roettiger et al., 1993; Pinkney et al., 1994; Burns, 1998; Sakelliou & Merrifield, 2000; Burns et al., 2002) where the bulk scale motions of the ICM are responsible for the ram pressure that bends the tails. And even though there is still debate over the origin of hybrid morphology radio sources (intrinsic or environmental), an asymmetric environment has been posited as one origin (Gopal-Krishna & Wiita, 2000; Kapińska et al., 2017) that could be created in a clustercluster merger (Gawroński et al., 2006). Therefore, it is not unexpected that a merging system could cause a change in morphology of jetted systems as we see in this work.
@@
We note that we only see HTs in merging systems (see tables in Appendix A). HTs are extreme versions of BT systems and are a product of stronger ram pressure from either a denser medium or a higher velocity of the source (; Miley et al. 1972; Terni de Gregory et al. 2017). Because only the most disturbed systems (HTs) are found in merging systems, this may indicate that merging systems allow for more opportunities for extreme ram pressure. An interesting follow-up would be to investigate the placement of these HT systems with respect to the cluster center.
@@
Our findings suggest that mergers have a marginally significant increased prevalence of radio AGN without an optical-AGN counterpart. If we assume that AGNRad can be roughly equated to an inefficiently accreting engine or low-excitation radio galaxies (LERGs) and the AGNRad-Opt can be roughly equated to an efficiently accreting engine or high-excitation radio galaxies (HERGs Hardcastle et al. (2009); Buttiglione et al. (2010); Best & Heckman (2012)), then we see that there are more LERGs than HERGs in merging clusters. Though we do not do a full HERG/LERG classification in this analysis (this was not the goal of this work, so potential biases exist such as the fact that this sample does not have an S/N requirement for the spectral lines needed to classify HERG/LERGs), we note that it would be interesting future work. In fact, the environment affecting the incidence of LERGs versus HERGs is not unprecedented. A few early works find that in the local Universe LERGs occupy more dense environments than HERGs (Prestage & Peacock, 1988; Allington-Smith et al., 1993; Zirbel, 1997; Best, 2004), which was confirmed by Miraghaei & Best (2017). Similarly, Ineson et al. (2015) also find a difference between the environments of HERGs and LERGs. On the other hand, Massaro et al. (2019, 2020) do not find a difference in HERG/LERG environments in the local Universe and note that the difference in environments of HERGs and LERGs seen in other works could be due to selection bias.
@@ 5.1 Discussion of Potential Sources of Uncertainty
There are several sources of error that we discuss here. An inherent source of error exists in the limitations of our classifications. Each source presents a unique radio emission profile, and while we aimed to encompass the bulk trends with the classification schemes, some intricacies may have eluded our classification process. Our morphological classification system does not encompass every type of radio galaxy morphology and was not designed to. We acknowledge the presence of jellyfish galaxies in our sample. Notably, we find examples of WC sources exhibiting faint extended emission that do not align exactly with the CE as defined here in this work but are noteworthy. In fact, two WC sources in A2255 are found in the existing literature to be ram-pressure-stripped galaxies (Ignesti et al., 2023). Further, we cross-match our sample with the LoTSS jellyfish catalog from Roberts et al. (2021) and find that four of our radio AGN are identified as jellyfish galaxies. We find that the jellyfish galaxies are equally spread between the WC, SC, CE, and E categories. However, we note that the LoTSS jellyfish catalog (Roberts et al., 2021) only includes data from LoTSS DR2 and that up to half of clusters in this work have LoTSS data that have been taken since DR2. Similarly, we identified sources displaying FR-type morphologies that do not precisely fit FRM or FR I/ II categories, potentially due to image artifacts, yet they exhibit distinctive and intriguing morphologies. To maintain the scope of this project, we limit our results to only the original Zooniverse classification results.
@@
An additional source of error that arises inherently in our classification-based project is disagreement. First, there are only 71 sources with complete agreement from all classifiers (consensus level 5), and the remaining 90 sources do not have complete agreement. On the other end of the spectrum within our data set, there were 26 sources that did not have a majority agreement (consensus level 2). Involving multiple classifiers was a necessary step to obtain robust classifications, even though it introduces some margin for error. There is a growing movement to move toward having multiple classifications or “tags” per source in the radio classification community in order to better describe the complex nature of the morphology of radio galaxies (Rudnick, 2021). For this exploratory work, we chose to have only one classification per object, but that leaves room for “error” in that we may not capture the full, complex morphology of the source.
@@
Another source of error is that of the dynamical state classification of the host cluster. It is difficult to obtain a binary classification of the dynamical state of clusters into “merging” and “relaxed.” Galaxy clusters are a direct result of hierarchical merging systems evolving over cosmic time, and the clusters we examine are snapshots in time, each characterized by varying degrees of merging activity. There is an ongoing debate surrounding the optimal approach for determining the dynamical state of a galaxy cluster. Our study is based on the Bilton et al. (2020) data set, which classifies clusters as merging or nonmerging based on a kinematic analysis. However, existing literature suggests that certain clusters that were classified as a ‘nonmerging’ cluster by Bilton et al. (2020) may in fact be undergoing merging processes based on X-ray observations that trace ICM turbulence (Nulsen et al., 2013; Blanton et al., 2011; Drabent et al., 2015; Yu et al., 2016; Wen & Han, 2013). Another way to classify the dynamical state of a cluster is based on the presence of central radio halos and radio relics, which can serve as an effective indicator of ongoing merger events. Notably, at least two clusters (A1367 and A1656) in this sample, classified as nonmerging using the Bilton et al. (2020) method, are identified as mergers based on radio halo observations (Kim et al., 1990; Gavazzi, 1978; Bonafede et al., 2022; Ge et al., 2019; Zhang et al., 2023).
@@
With these facts in mind, the clusters that Bilton et al. (2020) identified as merging and nonmerging via kinematics might not have that classification if one were to use a different type of analysis. Though it is difficult to get uniform, multiwavelength data for a large sample of clusters, and depending on the science goals, a multiwavelength approach would be the most robust best way to determine the dynamical state of a cluster.
@@
A last source of error is the choice to use a global RMS for the cutout contours. The choice of a global RMS per cluster inherently may lead to the loss of smaller-scale emission for individual sources where the RMS may be different from the global RMS. Given our original dataset from the Bilton et al. (2020) catalog, which comprises over 1000 sources, we were confronted with a trade-off between efficiency and sensitivity. Consequently, emission of lower surface brightness, such as that associated with ram pressure stripping, may not be fully captured. Existing work suggests at least two WC sources from A2255 are ram-pressure-stripped galaxies (Ignesti et al., 2023). However, in §3.1 we investigate the effect of choosing global versus a local RMS and find that our choice of a global, cluster RMS does not affect our results.